# Understanding Ethereum Mempool Security under Asymmetric DoS by Symbolized Stateful Fuzzing

### v3

@@ 6.4.2 Attack Evaluation on Reth
We report the total fees of normal transactions included in the blocks – the lower the fees are, the more successful the locking attack is. We also report the adversarial transaction fees as the attack cost shown in Figure 5. When the locking attack begins at the $45$-th block, the normal transaction fees increase because of the empty mempool on the Reth node which accepts some normal transactions in addition to the adversarial locking transactions. Then, after three blocks, the normal transaction fees quickly drop to near-zero Ether, showing the success of locking attacks. Interestingly, unlike the eviction attacks that cause low Gas utilization per block, the locking attacks just reduce the total Ether fees without reducing Gas utilization (i.e., the blocks produced under locking attacks from height $45$ to $60$ use the Gas of almost $100\%$ block limit). Due to the high Gas utilization, block $BasePrice$ keeps increasing during the locking attack. When reaching the $60$-th block, the $BasePrice$ is higher than the maximal transaction price tolerable by the attacker. In our experiment, the locking attacker stops at $60$-th block, and the normal transaction fees immediately recover to the “normal” level. In practice, the attacker can repeat sending eviction attacks to reduce the $BasePrice$ before mounting the locking attack again.
7 Evaluation of mpfuzz
7.1 Performance
Preliminary of GoFuzz: Upon each fuzzing iteration, GoFuzz [16] generates a bit-string of variable length, feeds it to our fuzzer code, and receives from our fuzzer a binary value of feedback, be either positive or negative. The GoFuzz further checks if this run increases the code coverage. If the feedback is positive and code coverage increases, GoFuzz adds the current bit-string to the seeds before running into the next iteration.
Baseline B1: Stateless fuzzer. We implement a stateless fuzzer in the GoFuzz framework described above. 1) Given a bit string generated by GoFuzz, our fuzzer code parses it into a sequence of transactions; the length of the sequence depends on the length of the bit string. It skips a bitstring that is too short. 2) The fuzzer sets up and initializes a Geth txpool instance with $m$ normal transactions, each sent from a distinct account, of nonce $1$, of value $1$ Ether (lower than the account balance), and of GasPrice $3$ wei. 3) The fuzzer sends the transaction sequence to the initialized mempool for execution. 4) It checks the mempool state after execution: If the test oracle is met, it emits the current transaction sequence as an exploit.
Baseline B2: Concrete-state-coverage fuzzer. We build the second baseline, a mempool fuzzer that takes concrete state coverage as feedback. Specifically, in each iteration, the fuzzer appends a new transaction to the current transaction sequence. Given the $m$-slot mempool, the fuzzer tries $m$ values for senders, nonces, GasPrice, and Ether amount. After sending the current transaction sequence to the mempool, it sorts the transactions in the mempool by senders and nonces. It then hashes the sorted transactions in the mempool and checks the presence of the hash digest in the seeds. If no hash exists, the current transaction sequence would increase the concrete-state coverage and is inserted to the corpus. The corpus is a FIFO queue where the seed inserted earlier has high priority to be de-queued.
Baseline B3: Invalid-tx energy fuzzer. Baseline B3 is similar to B2 except that it uses as the energy the number of invalid transactions on a mempool state. The seed with more invalid transactions in the mempool has higher energy or priority to be selected for the next round of fuzzing.
Baseline B4: feedback of no promising states: We design Baseline B4 for an ablation study on the effectiveness of using promising states as feedback in improving fuzzing performance. B4 is the same with mpfuzz except that it removes the state promising-ness from its feedback. Specifically, the feedback formula in B4 includes only state coverage ($st\_coverage$) and excludes state promisingness ($st\_promising$ as in Equation 7).
We implement Baseline B1 in Go/GoFuzz and implement B2, B3 and B4 in Python. We use a test oracle that can detect DETER attacks, that is, the number of invalid transactions in a mempool state equal $m-1$.
Table 4: Fuzzing Geth $v1.11.3$’s mempool (in minutes) by different approaches to detect Exploit $XT_{3}$.
Settings B1 B2 B3 B4 mpfuzz
$6$slot-$2$h Timeout $54$ $8$ $1.22$ $0.03$
$16$slot-$16$h Timeout Timeout $447$ Timeout $0.06$
Results: Table 4 reports the results under the small and medium MUT settings. For the small MUT, B1 cannot find any DETER attack in two hours. B2, B3 and B4 can find $XT_{3}$ in 54 minutes, $8$ minutes and $1.22$ minutes respectively. By contrast, mpfuzz finds the same exploit in $0.03$ minutes.
For the medium setting, with $16$ slots, baselines B1, B2, and B4 cannot find any DETER exploit in $16$ hours, and B3 can find Exploit $XT_{3}$ in $7.4$ hours. By contrast, mpfuzz finds the same exploit in under a minute.
The evaluation on different clients is in Appendix C.
7.2 True/False Positive Rates
This section evaluates the true/false positives of the short exploits found by mpfuzz. Recall that mpfuzz runs against a small mempool (MUT) and may introduce false positives, that is, the exploits that satisfy bug oracles in MUT but do not in the actual larger mempool. We formally define true and false positives as follows:
Definition 7.1 (TP/FP exploits)
Consider a short exploit found by mpfuzz on a MUT, $\langle{}st_{0},dc_{0}\rangle{},ops\Rightarrow{}\langle{}st_{n},dc_{n}\rangle{}$, and a (manually) extended exploit on an actual mempool, $\langle{}st_{0}^{\prime},dc_{0}^{\prime},ops^{\prime}\Rightarrow{}\langle{}st_{n}^{\prime},dc_{n}^{\prime}\rangle{}$.
The pair of exploits constitute a true positive eviction attack w.r.t. $\epsilon$ ($\lambda$), if.f. $st_{0}\cap{}st_{n}=\varnothing\land{}asym_{E}(st_{0},ops)<\epsilon\land{}\ st_{0}^{\prime}\cap{}st_{n}^{\prime}=\varnothing\land{}asym_{E}(st_{0}^{\prime},ops^{\prime})<\epsilon$
The pair of exploits constitute a false positive eviction attack w.r.t. $\epsilon$ ($\lambda$), if.f. $st_{0}\cap{}st_{n}=\varnothing\land{}asym_{E}(st_{0},ops)<\epsilon\land{}\ (st_{0}^{\prime}\cap{}st_{n}^{\prime}\neq{}\varnothing\lor{}asym_{E}(st_{0}^{\prime},ops^{\prime})>\epsilon)$.
True/false positives on locking attacks can be similarly defined from bug-oracle definitions 4.2 and 4.3.
We evaluate the false positives of the exploits found by mpfuzz under varying $\epsilon$ and $\lambda$. Specifically, we design a two-step experiment: First, given a MUT, we run mpfuzz with varying $\epsilon$ (and $\lambda$). Increasing $\epsilon$ allows mpfuzz to find more short exploits; for each distinct short exploit, we manually extend it to an actual exploit. Second, we evaluate the extended exploits on an actual mempool (in the same local experimental setting as described in § 9).
Table 5: True/false positives (TP/FP) of eviction and locking attacks discovered with varying $\epsilon$ and $\lambda$. ✓means satisfied.
Exploit $\epsilon$ $m=16$ (MUT) Default $m$ TP?
Eq. 1 $asym_{E}$ Eq. 1 $asym_{E}^{\prime}$
Geth-$XT_{1}$ $10^{-4}$ 00 00
Geth-$XT_{3}$09.09.09.09 $0.083$ $0.0002$
Geth-$XT_{6}$125.125.125.125 $0.125$ $0.0003$
Geth-$XT_{2}$ $0.2$ $0.167$ $0.0698$
Geth-$XT_{4}$ $0.23$ $0.208$ $0.0768$
Geth-$XT_{5}$ $0.23$ $0.208$ $0.0768$
Nethermind-$XT_{1}$ $10^{-4}$ 00 00
Nethermind-$XT_{4}$ $0.11$ $0.104$ $0.0008$
Nethermind-$XT_{7}$ $0.36$ $0.355$ $0.0012$
Besu-$XT_{2}$ $0.2$167.167.167.167 $0.0754$
Besu-$XT_{4}$ $0.23$208.208.208.208 $0.0083$
Erigon-$XT_{4}$ $0.23$ $0.208$ $0.0894$
Exploit $\lambda$ $m=16$ (MUT) Default $m$ TP?
Eq. 4 $asym_{D}$ Eq. 4 $asym_{D}^{\prime}$
Reth-$XT_{8}$ $0.34$ $0.34$ $0.015$
OpenEthereum-$XT_{9}$ $0.46$ $0.46$ $0.0439$
Table 5 presents the details of exploits discovered by mpfuzz. For instance, when setting $epsilon$ at $0.0001$ to find eviction attacks on Geth, mpfuzz discovered one short exploit $XT_{1}$ with $asym_{E}=0$ on MUT, which can be extended to an actual exploit on Geth mempool (of $m=6144$) with $asym_{E}^{\prime}=0<0.0001=\epsilon$. This makes it a true positive (marked by ✓in the table). Increasing $\epsilon$ to $0.23$ leads to discovering all six exploits $XT_{1-6}$ on Geth. All these exploits are true positives. The table also includes the results of eviction exploits found on other clients and locking exploits. Overall, the true-positive rate of mpfuzz remains at $100\%$ for $\epsilon\leq{}0.36$ and for $\lambda\leq{}0.46$ across Ethereum clients.
8 Discussions
Root causes of the exploits: We attribute all found exploits $XT_{1-9}$ to four distinct causes: 1) the presence of certain admission patterns (including $XT_{1}/XT_{2}/XT_{3}/XT_{4}/XT_{6}/XT_{7}$). For instance, $XT_{4}$ allows transaction replacement to cause latent overdrafts in place of valid transactions. 2) the absence of certain admission patterns to cause mempool locking (including $XT_{9}$ and $XT_{8}$). For instance, the cause of $XT_{8}$ on Reth is that Reth mempool is a FIFO queue, and it disallows transaction eviction of any kind, which is risky. 3) Inconsistency across multiple admission patterns, notably $XT_{7}$. In $XT_{7}$, the cause is that a mempool allows an eviction and its reversed eviction at the same time. 4) Evadable conditions to context-sensitive admission patterns, including $XT_{3}$ and $XT_{6}$. Both exploits work on Geth, where the mempool triggers the limit of transactions per send under an evadable condition (when the mempool stores more than $py_{3}^{\prime}=5120$ pending transactions).
Mitigation: We propose schemes to mitigate all newly discovered exploits comprehensively. In principle, given an exploit under a known cause, our mitigation design is to negate the cause. For instance, to mitigate $XT_{4}$, we propose declining transaction replacements that cause latent overdrafts. We have implemented this mitigation strategy on Geth, which was merged into the release of Geth $v1.11.4$ [8] . Similarly, to mitigate $XT_{7}$, we propose ensuring the consistency between eviction events and reversed events (e.g., if admitting $tx_{1}$ by evicting $tx_{2}$ is allowed, then admitting $tx_{2}$ by evicting $tx_{1}$ is not allowed). $XT_{1}/XT_{2}$ are mitigated by disallowing eviction of valid transactions by future/latent-overdraft transactions. $XT_{3}/XT_{6}$ are mitigated by removing the triggering condition of transaction sender limit (i.e., making $py_{3}^{\prime}=0$). $XT_{8}$ can be mitigated by re-enabling transaction evictions in certain cases, such as price-based transaction eviction (i.e., a pending transaction arriving later and of higher price can evict another pending transaction of lower price and admitted earlier).
If an exploit relies on multiple causes (e.g., $XT_{3}$ on Causes 1 and 4), negating one cause suffices to mitigate the exploit.
There are more sophisticated cases where one exploit relies on the presence of an admission pattern, and another exploit relies on the absence of the pattern. For instance, $XT_{5}$ relies on the presence of an admission policy named $SAP_{5}$ that turns child transactions into future ones (as in Geth), and the success of $XT_{9}$ relies on the absence of the same policy, that is, $\neg{}SAP_{5}$ (as in OpenEthereum).
To mitigate both $XT_{5}$ and $XT_{9}$, we propose a defense strategy in which the admission decision regarding $SAP_{5}$ is non-deterministic or randomized so that $SAP$ is not always on or off, making it hard for either $XT_{5}$ or $XT_{9}$ to always succeed.
[…]
Ethical concerns: When evaluating attacks, we only mounted attacks on the testnet and did so with minimal impacts on the tested network. For instance, our attack lasts a short period of time, say no more than $4$ blocks produced. We did not test our attack on the Ethereum mainnet. We also mask part of the block numbers in the screenshots (e.g., Figure 3) to prevent detailed attack inspection and reproduction. When reporting bugs, we disclose to the developers the mitigation design tradeoff and the risk of fixing one attack by enabling another attack (described above). To prevent introducing new bugs, we did not suggest fixes against turning-based locking ($XT_{5}$, $XT_{6}$) and locking ($XT_{8}$).
9 Conclusion and Future Works
Conclusion: This paper presents mpfuzz, the first mempool fuzzer to find asymmetric DoS bugs by exploring symbolized mempool states and optimistically estimating the promisingness of an intermediate state in reaching bug oracles. Running mpfuzz on popular Ethereum clients discovers new mempool-DoS vulnerabilities, which exhibit various sophisticated patterns, including stealthy mempool eviction and mempool locking.
Limitations and future works: The exploit generation in this work is not fully automated. Manual tasks include mempool reduction, mpfuzz setup (configuring $\epsilon$ and symbols), exploit extension, etc. Automating these tasks is the future work.
mpfuzz targets a victim mempool of limited size: the vulnerabilities found in this work are related to transaction eviction, which does not occur in a mempool of infinite capacity. Thus, the mpfuzz workflow cannot find exploits in the mempool of infinite capacity. It is an open problem whether a mempool of large or infinite capacity has DoS vulnerabilities.
10 Acknowledgments
The authors thank USENIX Security reviewers and Kangjie Lu for the discussion in the early stage of this work. Yibo Wang, Wanning Ding, Zhihua Yang, and Yuzhe Tang were partially supported by two Ethereum Foundation academic grants and NSF grants CNS-2139801, CNS-1815814, DGE-2104532. Kai Li is partially supported by NSF grant CNS-2347486 and one Ethereum academic grant.
Appendix A Observations Motivating mpfuzz Design
(In)feasibility of code-coverage-only stateless fuzzing: To fuzz a mempool, a baseline design is to use a code-coverage-based fuzzer. In each iteration, the fuzzer generates a bit string, parses it into a sequence of transactions and sends them to the tested mempool for execution. Upon each end state, it checks whether the test oracle defined above is satisfied. This design takes code coverage as the only feedback and is stateless because the feedback does not consider the end state.
The stateless design would be ineffective in fuzzing mempool or finding ADAMS exploits. For example, consider fuzzing a three-slot mempool to find a DETER-X exploit consisting of three future transactions [28]. Suppose the current mempool input is a sequence of one future transaction. After mutation, it may try an input of two future transactions, which, however, does not increase code coverage (the second future transaction is executed by the same code path in a mempool as the first future transaction). Thus, the fuzzer discards the two-future-transaction input and misses finding the three-future-transaction exploit. We validate the ineffectiveness of stateless fuzzer by experiments in § 7.
A more promising design is stateful mempool fuzzing, in which the mempool state is included in the feedback to guide next-iteration fuzzing. In the previous example of finding a three-future-transaction exploit (DETER-X), sending two future transactions leads to a different mempool state than sending one. A simple state-coverage-based fuzzer would view this as positive feedback and further explore this direction toward finding the three-future-transaction exploit.
(In)feasibility of existing consensus fuzzers: Existing consensus fuzzers [25, 29, 36] cannot detect mempool DoS or ADAMS bugs. Specifically, Loki’s test oracle is detecting system crashes, while a successful ADAMS attack does not necessarily trigger system crashes. Using Loki to detect ADAMS would cause many false negatives.
Fluffy detects the unsynchronizable difference of post-consensus states as vulnerability. However, the mempools in two clients, say Geth and Besu under benign transactions, would be different. In other words, post-consensus state differences across clients could indicate insecurity, but the difference in pre-consensus states across clients, like mempool, does not mean insecurity. Using Fluffly to detect ADAMS could cause false positives.
Tyr’s test oracle detects property violations on post-consensus states. Particularly, while Tyr models state liveness, their definition “valid transactions should be executed eventually” is, unfortunately, over-simplifying and fails to model the legitimate pre-consensus cases in real Ethereum clients; for instance, valid transactions of low price can be dropped by Ethereum mempools and will not be eventually executed. Besides, they model double-spending transactions as the only type of invalid transaction. In Ethereum, double-spending transactions are of the same nonce. And there are many more sophisticated invalid transactions in Ethereum that Tyr does not model, including future transactions, latent overdrafts, etc.
We analyze the case of existing consensus fuzzers in detecting known DETER attacks [28]. Loki cannot detect DETER because DETER does not cause system crashes. Fluffy does not model future transactions and can not find the DETER attacks that rely on future transactions. Similarly, Tyr that does not model sophisticated invalid transactions would miss detecting DETER attacks.
A.1 Further Rationale of Symbolization
We now describe the design rationale of our symbolization technique.
  • The first observation is that real-world mempools admit transactions based on the symbolic value, not concrete value, of nonce and Ether amount. Our design reflects this intuition. For instance, symbols of valid transactions, e.g., $\mathcal{P},\mathcal{C},\mathcal{N}$, are instantiated by mpfuzz to a minimal amount $1$ Ether. As long as the transferred value does not exceed the sender balance, transactions remain valid (i.e., no overdraft). Admission decisions remain the same for two valid transactions despite their difference in nonces or values.
    Likewise, Symbol $\mathcal{F}$ is instantiated by mpfuzz to fixed nonce $n+1$: As long as the nonce is non-consecutive, no matter what specific value it takes (e.g., $3$ or $m+1$), the mempool would deem it as a future transaction and makes the same admission decision.
    As a side note, the nonce of Symbol $\mathcal{F}$ is fixed at a large nonce $n+1$ (we denote it by far future transactions) instead of a small nonce, say $5$ (we denote it by near future transactions), because near future transactions’ nonce could be reconnected as consecutive and lead to duplicated states.
  • The second observation is that restricting eviction/replacement victims to only parent transactions can facilitate finding exploits quickly. Specifically, while any transactions can be evicted/replaced, mpfuzz evicts/replaces only parent transactions; such a strategy can cause maximal damage (i.e., the most child transactions turned) while saving the search space. Our symbol design reflects this idea. For instance, Symbol $\mathcal{R}$ covers any transaction of the same sender and nonce as a transaction in state $st$, but it is instantiated by mpfuzz to only the transaction of nonce $1$ (i.e., replacing a parent transaction).
Appendix B Case Study: How mpfuzz Finds Exploits
We describe how mpfuzz finds an exploit by presenting a case study on finding $XT_{6}$ in the latest Geth $\geq{}v1.11.4$.
Mempool reduction: Recall that a Geth mempool has a capacity of $m^{\prime}=6144$ slots, and its transaction-admission policies are characterized by three essential parameters: admitting up to $py_{1}^{\prime}=1024$ future transactions and limiting up to $py_{2}^{\prime}=16$ pending transactions from any senders when more than $py_{3}^{\prime}=5120$ pending transactions are residing in the mempool. We set up the MUT to run the same codebase or the same admission policy but with smaller parameters: $m=3,py_{1}=1,py_{2}=2,py_{3}=2$ What follows is a description of how mpfuzz find a short exploit on this MUT.
Fuzzing: How mpfuzz automatically finds exploits: Initially, the mempool is filled with $m=3$ normal transactions. That is, the initial symbolized state is $\mathcal{NNN}$. The seed corpus $sdb$ initially contains an empty string. mpfuzz retrieves the empty string and appends to it with different symbols $\mathcal{O}_{0}$, $\mathcal{C}_{0}$, $\mathcal{P}_{0}$. Because of the initial state $\mathcal{NNN}$, only Symbol $\mathcal{P}_{0}$ is feasible. It generates the mutated input $\mathcal{P}$ and instantiates it to a parent transaction of a higher GasPrice than normal transactions (as described in § 5). Sending the input to the mempool gets the transaction admitted, leading to transitioned state $st_{1}=\mathcal{NNP}$. This is a new state that is not in the corpus, and it evicts more normal transactions $\mathcal{N}$ than the previous state $st_{0}$; the input produces positive feedback, and the associated input-state pair $\langle{}P,st_{1}=\mathcal{NNP}\rangle{}$ is added to the corpus.
Next, mpfuzz retrieves from $sdb$ a seed by high energy. According to Table 2, the energy of state $\mathcal{NNN}$ is $\frac{1}{3*3}*0$, and the energy of state $\mathcal{NNP}$ is $\frac{1}{3*2+4}*1=1/10>0$. Hence, seed $\mathcal{NNP}$ is selected. mpfuzz tries input mutation and appends to the selected input $\mathcal{P}$ one of four new symbols, that is, $\mathcal{O}_{1}$, $\mathcal{C}_{1}$, $\mathcal{P}_{0}$ or $\mathcal{P}_{1}$. On state $\mathcal{NNP}$, 1) mutation transaction instantiated from symbol $\mathcal{O}_{1}$ is declined admission. 2) Mutation transaction from $\mathcal{C}_{1}$ is admitted, transitioning the state to $\mathcal{NPC}$, which produces positive feedback and is added to $sdb$. 3) Likewise, transaction $\mathcal{P}_{0}$ is admitted and produces state $\mathcal{NPP}$ of positive feedback; the mutated input is also added to $sdb$. 4) Mutation $\mathcal{P}_{1}$ is admitted but produces an identical state with mutation $\mathcal{P}_{0}$; thus, the mutated input is not added to $sdb$.
Now, there are four seeds in $sdb$: $\mathcal{NNN}(0)$, $\mathcal{NNP}(0)$, $\mathcal{NPC}(1/8)$, and $\mathcal{NPP}(1/12)$. In parentheses are their energy numbers. The seed of the highest energy $\mathcal{NPC}$ is selected ( 1 in Figure 6). mpfuzz then mutates $\mathcal{NPC}$ with $4$ possibilities, that is, $\mathcal{O}_{1}$, $\mathcal{C}_{1}$, $\mathcal{P}_{0}$, and $\mathcal{P}_{1}$, which produce two end states with positive-feedback, that is, $\mathcal{PPC}$ and $\mathcal{PCP}$. They are added to the $sdb$ with energy $\mathcal{PPC}(1/10)$ and $\mathcal{PCP}(1/10)$.
Let’s say $\mathcal{PPC}(1/10)$ is selected ( 2). mpfuzz mutates input $\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}$ with $7$ mutation transactions, that is, $\mathcal{O}_{1}$, $\mathcal{O}_{2}$, $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{P}_{0}$, $\mathcal{P}_{1}$, $\mathcal{P}_{2}$, which produces one end state with positive feedback, that is, mutation $\mathcal{C}_{1}$ is admitted and transitions state to $\mathcal{FPC}$. After that, $\mathcal{FPC}$ is the one with the highest energy and is chosen for the next-round fuzzing ( 4). mpfuzz mutates input $\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}\mathcal{C}_{1}$ with $4$ mutations and produces $2$ state transitions with positive feedback. Upon state $\mathcal{FPC}$, mutation $\mathcal{C}_{1}$ is admitted and leads to state $\mathcal{FEE}$, which satisfies the bug oracle of eviction attacks under $\epsilon=0$. The algorithm then emits the found short exploit: $\langle{}st_{0}=\mathcal{NNN},dc_{0}=\varnothing\rangle{},ops=\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}\mathcal{C}_{1}\mathcal{C}_{1}$ (recall Definition 4.1). The snapshot of the state-search tree is depicted in Figure 6.
Exploit extension: Given the short exploit automatically found on MUT (with $m=3,py_{1}=1,py_{2}=2,py_{3}=2$), the next step is to extend it to a longer exploit functional on the actual Geth mempool (with $m^{\prime}=6144,py_{1}^{\prime}=1024,py_{2}^{\prime}=16,py_{3}^{\prime}=5120$).
Exploit extension requires manual efforts: After identifying exploit $\langle{}st_{0}=\mathcal{NNN},dc_{0}=\varnothing\rangle{},ops=\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}\mathcal{C}_{1}\mathcal{C}_{1}$ is unique, we extend it to the longer exploit by ensuring the same admission event occurs on the actual mempool as on the smaller MUT. In this process, it tries the next transaction of the same sender with the previous one but with an incremented nonce. If it fails, it switches to the next sender. It also tries transaction fees/prices based on measuring the fees of actual normal transactions, which may not be fixed as $3$ in the mpfuzz setting.
Appendix C Additional Eval. of mpfuzz Performance
We further run our mpfuzz on different clients, including Geth $v1.10.11$, Geth $v1.11.4$, Erigon $v2.42.0$, and Nethermind $v1.18.0$. The tested mempool is configured at $16$ slots. We report the number of exploits found by mpfuzz in a $16$-hour period. Figure 8 presents the result that mpfuzz finds $4096$ short exploits in the first $8$ hours on Geth-$v1.10.11$. On Geth-$v1.11.4$, mpfuzz quickly finds one exploit (i.e., exploit $XT_{6}$ as described in § 6.1) within $2$ minutes and does not find anymore in the next $16$ hours. On Nethermind, it finds one exploit (exploit $XT_{1}$) within two minutes and finds the next one (exploit $XT_{4}$) near the end of $16$-th hour. On Erigon, it finds one exploit ($XT_{4}$) in the $16$-th hour.
Figure 7: # exploits found
Figure 8: # states covered
Figure 8 shows the number of states that are explored by mpfuzz on a $16$-slot mempool in 16 hours. mpfuzz explores $749$ and $91428$ states on Geth $v1.10.11$ in $2$ minutes and $16$ hours respectively. On Geth $v1.11.4$, mpfuzz explores $372$ and $23286$ states in $2$ minutes and $16$ hours respectively. However, the performance of Erigon and Nethermind is much lower than that of Geth. mpfuzz explores $12$ and $1239$ states on Erigon in $2$ minutes and $16$ hours respectively. On Nethermind, mpfuzz explores $15$ and $1433$ states in $2$ minutes and $16$ hours, respectively. The reason mpfuzz is more performant on Geth is that we implemented an external API on Geth to initialize the mempool; in each iteration of fuzzing on Geth, mpfuzz calls the API to initialize the mempool. In contrast, on Nethermind and Erigon, mpfuzz restarts the client in each iteration, which is consuming.
Appendix D Additional Attack Evaluation
D.1 Experiment on a Single Victim Node

### v4

@@ 7-B2 Attack Evaluation on Reth
We report the total fees of normal transactions included in the blocks – the lower the fees are, the more successful the locking attack is. We also report the adversarial transaction fees as the attack cost shown in Figure 5. When the locking attack begins at the $45$-th block, the normal transaction fees increase because of the empty mempool on the Reth node which accepts some normal transactions in addition to the adversarial locking transactions. Then, after three blocks, the normal transaction fees quickly drop to near-zero Ether, showing the success of locking attacks. Interestingly, unlike the eviction attacks that cause low Gas utilization per block, the locking attacks just reduce the total Ether fees without reducing Gas utilization (i.e., the blocks produced under locking attacks from height $45$ to $60$ use the Gas of almost $100\%$ block limit as shown in Figure 6). Due to the high Gas utilization, block base fees keep increasing during the locking attack. When reaching the $60$-th block, the base fee is higher than the maximal transaction price tolerable by the attacker. In our experiment, the locking attacker stops at $60$-th block, and the normal transaction fees immediately recover to the “normal” level. In practice, the attacker can repeat sending eviction attacks to reduce the base fee before mounting the locking attack again.
Figure 6: Gas usage of attacks $XT_{8a}$
Figure 6 shows the Gas usage of attack transactions and normal transactions in $XT_{8a}$. During the locking attack from block height $45$-th to $60$-th, the attacker sends locking transactions to occupy all the empty slots. It makes the Gas usage of the attacker almost 100% block limit while the Gas usage of the normal transactions drops to around 00. When empty slots are created by arriving blocks, normal transactions compete with the attack transactions to admit into the empty slots. Thus, some normal transactions are included in blocks during the locking attack.
We also manually designed another locking variant for Reth. Suppose the Reth node does not run block validation, and its mempool is initially synchronized with the mempool of a block validator. The attacker first evicts $10,000$ normal transactions from the validator’s mempool. Then, the Reth node’s mempool is permanently locked by the $10,000$ normal transactions evicted from the validator mempool. Because these 10,000 transactions will never be included in blocks, they would exist in the Reth node’s mempool forever, permanently locking it in its current state.
Figure 7: Evaluation of locking attack $XT_{8b}$ on Reth
Evaluation of $XT_{8b}$: To evaluate $XT_{8b}$, we initially set up the attacker node connected to the validator node running Geth, the workload generator node connected to the victim non-validator Reth node, and the Reth node connected to the validator. The system is run long enough to produce a certain number of blocks and the normal transactions are propagated to both validator and on-validator nodes’ mempools. We then mount the eviction attack ($XT_{6}$) directly to the validator Geth node that evicts the mempool transactions on the validator node and prevents them from being included in the blocks.
By mounting eviction attacks between the $30$-th to $45$-th blocks in Figure 7, the Reth mempool is successfully locked afterward: From the figure, after the $45$-th block, the total fees of normal transactions included in blocks remain at zero Ether per block until the experiment ends on the $60$-th block.
7-C Evaluation of Locking on Geth
Figure 8: Experimental setup for locking attacks on Geth/OpenEthereum
(a) Attack damage
(b) Attack cost
Figure 9: Mounting locking attacks on a Geth node in testnet.
Design rational: Recall that OpenEthereum adopts an overly strict policy that determines eviction by only considering the fees of childless transactions. While this policy is intended to prevent turning attacks ($XT_{5}$), it is vulnerable to locking attacks. Because OpenEthereum has been deprecated from the operational Ethereum 2.0 networks, we re-implement its admission policy on Geth $v1.11.4$ and run the turning-hardened Geth on a victim under locking attacks.
Settings: We set up an experimental platform for evaluating locking attacks on Goerli testnet. The victim node runs the turning-hardened Geth client (based on $v1.11.4$) described above. The attacker node runs an instrumented Geth client (based on $v1.11.4$) at the execution layer such that it can mount locking attacks. Specifically, the Geth instrumentation enables the attacker node to monitor the local mempool and send a crafted transaction to the victim node upon each new empty slot created (by block arrival). The network topology is depicted in Figure 8.
In the experiment, we first run our three nodes for $350$ minutes and then turn on the locking attack on the attacker node. We collect the benign transactions received by the attacker node (the number of which is denoted by $N_{a}$), the crafted transactions the attacker sends (the number denoted by $N_{c}$), and the benign transactions received by the measurement node (the number denoted by $N_{m}$). We use Metric $\frac{N_{a}}{N_{m}}$ to report the (inverse) of the attack success rate – Under the same transactions received by the measurement node, the fewer transactions the attacker node receives from the victim node, the more successful the victim mempool is locked.
Evaluation of locking attacks: We report in Figure 9a the locking attack’s success rate over the period of $350$ minutes. Each tick on the $X$ axis represents $1$ minute; that is, $N_{a}$/$N_{m}$ is the number of benign transactions received by the attacker/measurement node every $1$ minute. It is clear that the locking attack is highly successful – Right after the attack is launched at the $180$-th minute, the rate of normalized transactions received from the victim node drops essentially to zero. And it persists over the period where the attack lasts.
We also calculate the attack cost by the total fees of the attacker’s transactions included in the blockchain. In the experiment, we wait three hours after the attack stops and consider all the blocks produced until then. Not a single attacker’s transaction sent in our experiment is included in the blockchain. This may be due to the low fee with which crafted transactions are sent.
Understanding crafted transaction inclusion is nondeterministic and dependent on benign transactions, we estimate the worst-case attack cost, which is the total fees of crafted transactions sent by the attacker (i.e., when all the crafted transactions are included in the blockchain). Figure 9b reports the worst-case locking attack cost over time; on average, the attack, as successful as reported in Figure 9a, consumes $0.0004$ Ether per block. By comparison, it is $0.028$ Ether of transactions included per block, which is the minimal cost of the baseline transaction-spamming attack [2].
7-D Evaluation of Found Exploits on Single Victim Node